OpenStudy (anonymous):

To achieved the velocity that you just computed, what must have been the potential difference (V) through which the proton was accelerated?

OpenStudy (anonymous):

W = Ek K*E = 1/2*m*v^2, (from 1eV = 1.6*10^-19 J) 1.6*10^-19*E = 1/2*1.67*10^-27*(3.9*10^5)^2, (from velocity of previous exercise) E = 820 V